# Automated Knowledge Concept Annotation and Question Representation Learning for Knowledge Tracing

### v1

@@ 5.2 Ablation Studies
Quality of the automated KC annotation: Here, we perform an ablation for module 1 of our framework (in Sec. 3.1) where we assess the quality of the KC annotations. We compare against two ablations: (a) the original KC annotations from the dataset and (b) the annotations from our KCQRL but without leveraging the solution steps. For the ablation study, we ran our automated evaluation via a different LLM, i.e., Llama-3.1-405B (Dubey et al., 2024), to avoid the potential bias from using the same model as in our KC annotation.
Table 3 shows the pairwise comparison of three KC annotations (i. e., ours and two ablations) based on 5 criteria, where the LLM model is prompted to choose the best of the given two annotations for each criterion. The two key observations are: (1) Both annotations from KCQRL (i.e., with and without solution steps) are clearly preferred over the annotations from the original dataset, confirming the findings from Moore et al. (2024). (2) Overall, KC annotations from our full framework are preferred over the annotations without solution steps, confirming our motivation of our KC annotation. Further details and example KC annotations are in Appendix G. Takeaway: A large performance gain is due to module 1 where we ground the automated KC annotations in chain-of-thought reasoning.
Table 3: Ablation study showing the relevance of automated KC annotation. We report the quality (in %) for different KC annotations.
@@
Fig. 5 plots the embeddings of questions. For all five baselines, the representations of questions from the same KC are scattered very broadly. In comparison, our KCQRL is effective in bundling the questions from the same KC in close proximity as desired. Takeaway: This highlights the importance of our CL training in Sec. 3.2.
Fig. 6 shows the performance of KT models with these five baseline embeddings in comparison to the default implementation and our complete framework. In most cases, baseline embeddings perform better than the default implementation of KT models, which again highlights the importance of semantic representations (as compared to randomly initialized embeddings). Further, our KCQRL is consistently better than the baseline embeddings, which highlights the benefit of domain-specific representation learning. Takeaway: A large performance gain is from module 2, which allows our framework to capture the semantics of both questions and KCs.
Figure 5: Visualization of question embeddings. For better intuition, we chose the same question as in the example from Fig. 2, and, for each of its KCs, we color the question representations sharing the same KC. Evidently, our CL loss is highly effective.
@@
Ablations for CL training: To demonstrate the effectiveness of our CL loss, we implement the following ablations: (f) KCQRL w/o false negative elimination, whose CL loss is calculated without the blue indicator function in Eq. 2 and Eq. 4, and (g) KCQRL w/o sol. steps, whose CL loss is calculated based only on $\mathcal{L}_{\textrm{question}_{i}}$ by ignoring $\mathcal{L}_{\textrm{step}_{i}}$.
@@ 6 Related Work
Knowledge tracing (KT) aims to model the temporal dynamics of students’ interactions with learning content to predict their knowledge over time (Corbett & Anderson, 1994). A large number of machine learning models, including deep neural networks, have been proposed for this purpose (Piech et al., 2015; Yeung & Yeung, 2018; Lee & Yeung, 2019; Sonkar et al., 2020; Long et al., 2021; Liu et al., 2023a; Chen et al., 2023; Zhang et al., 2017; Yeung, 2019; Guo et al., 2021; Pandey & Karypis, 2019; Choi et al., 2020a; Ghosh et al., 2020; Liu et al., 2023b; Huang et al., 2023).
@@
KC annotation aims to infer the KCs of a given exercise in an automated manner. (i) One research line infers KCs from students’ learning histories as latent states (e.g., Barnes, 2005; Cen et al., 2006; Liu et al., 2019; Shi et al., 2023). However, these latent states lack interpretability, limiting their use in KT. (ii) Another approach reformulates KC annotation as a classification task (e.g., Patikorn et al., 2019; Tian et al., 2022; Li et al., 2024a; b), but this requires large annotated datasets, making it often infeasible. (iii) Recent works have also explored KC annotation in free-text form (Moore et al., 2024; Didolkar et al., 2024), but their annotations are not informed by the question’s solution steps, which we show is suboptimal in our ablations. In our work, we leverage LLMs’ reasoning abilities to generate step-by-step solutions and produce grounded KCs for each question.
@@ 7 Discussion
Implications: Our framework allows even the simplest KT models to reach near state-of-the-art (SoTA) performance, suggesting that much of the existing literature may have “overfit” to the paradigm of modeling sequences based only on IDs. Future KT research can advance by focusing on sequential models “designed to” inherently understand semantics. As a first step, we release an English version of the XES3G5M dataset with full annotations, including question content, solutions, and KCs.
@@ Appendix A Related Work (Extended)
Knowledge tracing (KT) aims to model the temporal dynamics of students’ interactions with learning content to predict their knowledge over time (Corbett & Anderson, 1994). Early works in KT primarily leveraged logistic and probabilistic models (Corbett & Anderson, 1994; Cen et al., 2006; Pavlik et al., 2009; Käser et al., 2017; Vie & Kashima, 2019). Later, many KT works focused on deep learning-based approaches, which can loosely be grouped into different categories: (1) deep sequential models that leverage auto-regressive architectures such as RNN or LSTM (Piech et al., 2015; Yeung & Yeung, 2018; Lee & Yeung, 2019; Liu et al., 2019; Nagatani et al., 2019; Sonkar et al., 2020; Guo et al., 2021; Long et al., 2021; Shen et al., 2022; Chen et al., 2023; Liu et al., 2023a); (2) memory augmented models to externally capture the latent states of the students during the learning process (Zhang et al., 2017; Abdelrahman & Wang, 2019; Yeung, 2019); (3) graph-based models that use graph neural networks to capture the interactions between knowledge concepts and questions (Nakagawa et al., 2019; Song et al., 2022; Cui et al., 2024); and (4) attention-based models that either employ a simple attention mechanism or transformer-based encoder-decoder architecture for modeling student interactions (Pandey & Karypis, 2019; Pandey & Srivastava, 2020; Choi et al., 2020a; Ghosh et al., 2020; Huang et al., 2023; Im et al., 2023; Liu et al., 2023b; Yin et al., 2023; Ke et al., 2024; Li et al., 2024c).
Outside these categories, Lee et al. (2024) fits entire history of students exercises into the context window of a language model to make predictions. Guo et al. (2021) designs an adversarial training to have robust representations of latent states, and Zhou et al. (2024) develops a generative model to track the knowledge states of the students. However, the existing works (1) require a comprehensive mapping between knowledge concepts and questions, and (2) overlook the semantics of questions’ content and KCs. Our KCQRL framework effectively addresses both limitations.
@@
KC annotation aims at inferring the KCs of a given exercise in an automated manner. (i) One line of research infers KCs of questions by learning patterns from students’ learning histories in the form of latent states (e.g., Barnes, 2005; Cen et al., 2006; Liu et al., 2019; Shi et al., 2023). Yet, such latent states lack interpretability, because of which their use in KT is typically prohibited. (ii) Another line of research reformulates KCs annotation as a classification task, typically via model training (Patikorn et al., 2019; Tian et al., 2022) or LLM prompting (Li et al., 2024a; b). Although these works do not require human experts at inference, they still require them to curate a large annotated dataset for the model to learn the task. (iii) Recent works also explored the KC annotation in free-text form (Moore et al., 2024; Didolkar et al., 2024). However, their KC annotations are not informed by the solution steps of the question. We show that this is suboptimal in our ablations.
@@
In our work, we leverage the reasoning abilities of LLMs (Wei et al., 2022; Fu et al., 2023; Shridhar et al., 2023; Wang et al., 2023; Yu et al., 2024) to generate the step-by-step solution steps of a given question and therefore generate the associated KCs in a grounded manner.
@@
Many works designed tailored CL losses (e.g., Oord et al., 2018; Chen et al., 2020; He et al., 2020; Ozyurt et al., 2023b) to improve the representations of textual data in various domains. Examples are information retrieval (Karpukhin et al., 2020; Khattab & Zaharia, 2020; Lee et al., 2023; Sun et al., 2023a; Zhu et al., 2023; Zhao et al., 2024), textual entailment (Gao et al., 2021; Li & Li, 2024; Ou & Xu, 2024), code representation learning (Zhang et al., 2024), and question answering (Yue et al., 2021; 2022). To our knowledge, we are the first to leverage a CL loss for the representation learning of questions specific to KT tasks.
@@ Appendix B Dataset Details
XES3G5M (Liu et al., 2023c): XES3G5M contains the history student exercises for 7,652 unique questions. These questions are mapped to 865 unique KCs in total. It has a total of 18,066 students’ learning histories and a total of 5,549,635 interactions for all students. Fig. 8(a) shows the distribution of interaction numbers across students. Of note, XES3G5M dataset is originally provided in Chinese, which we translated to English to make it compatible to our framework. Details are given below.
@@
Translation: The original XES3G5M dataset is provided in Chinese. To make it compatible with our KCQRL framework, we first translated the question contents to English. We did the translation via Google Translate via deep-translate library in Python.
@@
To make the fill-in-the-blank questions consistent in the dataset (and also consistent with our other dataset Eedi), We prompt GPT-4o to convert these questions to proper question phrases. For example, “…There are ___ different ways to wear it.” is converted to “…How many ways are there to wear it?”. We provide our prompt template in Appendix B.1.
Eedi (Eedi, 2024): Eedi contains the history student exercises for 4,019 unique questions. These questions are mapped to 1215 unique KCs in total. It has a total of 47,560 students’ learning histories and a total of 2,324,162 interactions for all students. Fig. 8(b) shows the distribution of interaction numbers across students.
@@ Appendix C KT Models Details
In our KCQRL framework, we report the performance of 15 KT algorithms on two real-world large online math learning datasets. The summary of these KT algorithms are given below:
  • DKT (Piech et al., 2015): It is the first deep learning-based KT algorithm. It uses the LSTM to model the temporality in students’ learning histories. The original DKT turns each KC into a one-hot or a random vector. The recent implementation from pykt (Liu et al., 2022b) learns the embeddings for each KC, which are then processed by an LSTM layer.
  • DKT+ (Yeung & Yeung, 2018): It extends the above DKT model by a custom regularization approach. Specifically, it adds a reconstruction loss to the last exercise’s prediction. Further, it adds a regularization term to make the predictions of the same KC consistent across the time dimension.
@@
  • IEKT (Long et al., 2021): It models the students’ learning histories via a RNN. In addition, it has two additional neural network modules, namely, for student cognition and knowledge acquisition estimation.
  • AT-DKT (Liu et al., 2023a): It has the same model backbone as in DKT. On top of it, AT-DKT addresses two auxiliary tasks: question tagging prediction and student’s individual prior knowledge prediction.
  • QIKT (Chen et al., 2023): It combines three neural network modules to model the learning processes: question-centric knowledge acquisition module via an LSTM, question-agnostic knowledge state module via another LSTM and question-centric problem solving module via a MLP.
@@
  • simpleKT (Liu et al., 2023b): As a simple but tough-to-beat baseline, this model employs and ordinary dot product as an attention mechanism to the embeddings of questions.
  • sparseKT (Huang et al., 2023): On top of simpleKT, this model further introduces a sparse attention mechanism such that, for the next exercise’s prediction, the model attends to at most $K$ exercises in the past.
@@ Appendix D Implementation Details
Here, we explain the implementation details of each part of our KCQRL framework.
@@
KC Annotation (Sec. 3.1): We leverage the reasoning abilities of OpenAI’s GPT-4o [^6] https://platform.openai.com/docs/models/gpt-4o model as our LLM $P_{\phi}(\cdot)$ at each step. We set its temperature parameter to 0 to get deterministic answers from the model. For each question in the dataset, the GPT-4o model is prompted three times in total: once for solution steps generation, once for KC annotation, and once for solution step-KC mapping. Overall, the cost is around 80 USD for XES3G5M dataset with 7,652 questions and 50 USD for Eedi dataset with 4,019 questions.
@@
Representation learning of questions (Sec. 3.2): We train BERT (Devlin, 2019) as our LLM encoder $E_{\psi}(\cdot)$. For the elimination of false negative pairs, we use HDBSCAN (Campello et al., 2013) as the clustering algorithm $\mathcal{A}(\cdot)$. The clustering is done over the Sentence-BERT (Reimers & Gurevych, 2019) embeddings of KCs, which is a common practice in identifying relevant textual documents (Liu et al., 2022a; Ozyurt et al., 2023a). For HDBSCAN clustering, we set the minimum cluster size to 2, the minimum samples to 2, and the metric to cosine similarity between the embeddings. For CL training, we train BERT (Devlin, 2019) as our LLM encoder. To better distinguish three three types of inputs, i. e. question content, solution step, and KC, we introduce three new tokens [Q], [S], [KC] to be learned during the training. These new tokens are added at the beginning of the question content, solution step, and KC, respectively, during both training and inference. For all input types, we use [CLS] token’s embeddings as the embeddings of the entire input text. For the training objective, we set $\alpha=1$. Since both losses $\mathcal{L}_{\textrm{question}_{i}}$ and $\mathcal{L}_{\textrm{step}_{i}}$ are already normalized by the number of KCs and steps, our initial exploration ($\alpha$ varying from 0.2 to 5.0) indicated that setting $\alpha=1$ yields the best performance. We train the encoder for 50 epochs with the following hyperparameters: batch size = 32, learning rate = 5e-5, dropout = 0.1, and temperature (of the similarity function) = 0.1. For training $E_{\psi}(\cdot)$, we use an Nvidia Tesla A100 with 40GB GPU memory. The entire training is completed under 4 hours.
@@
Improving KT via learned question embeddings: We adopt the following strategy to ensure a fair evaluation of KT models and their improvements via our novel KCQRL framework. We first did a grid search over the hyperparameters of each KT model (see Table D for parameters) to find the best configuration for each model. Then, we replaced the model’s embeddings with our learned question embeddings and trained/tested the KT models with the same configurations found earlier (i. e., no further grid search is applied). To ensure a fair evaluation between the KT model and in comparisons to the versions improved by our framework, we fixed their embedding dimensions to 300. As BERT embeddings’ default dimensionality is larger (i. e., 768), we just added a linear layer (no non-linear activation function is added) on top of the replaced embeddings to reduce its dimensionality to 300 to ensure that model capacities are subjected to a fair comparison. As KT models are much smaller than the LLM encoders, we now used NVIDIA GeForce RTX 3090 with 24GB GPU memory for even larger batch sizes.
@@
Evaluation of KC-centric KT models: Some KT models (such as DKT (Piech et al., 2015), simpleKT (Liu et al., 2023b), etc.) expand the sequence of questions into the sequence of KCs for both training and inference. For instance, if there are 3 questions with 2 KCs each, they transform the sequence $\{q_{1},\,q_{2},\,q_{3}\}$ into $\{c_{11},\,c_{12},\,c_{21},\,c_{22},\,c_{31},\,c_{32}\}$ and assign the labels of original questions to each of their corresponding KCs for training. This paradigm causes information leakage in the evaluation as highlighted by Liu et al. (2022b). The reason is that, during the prediction of $c_{32}$, the label of $c_{31}$ is already given as the history, which is the same label (to be predicted) for $c_{32}$. To eliminate the information leakage, we followed the literature (Liu et al., 2022b; 2023c) and applied the following procedure. Consider the above example where we predict the label of $q_{3}$. (1) We provided the expanded KC sequence of earlier questions as before, i. e., $\{c_{11},\,c_{12},\,c_{31},\,c_{32}\}$. (2) We next appended $c_{31}$ and $c_{32}$ separately to the given sequence and run the predictions independently. (3) Finally, we aggregated these predictions by taking their mean and used it as the model’s final prediction. Important note: With our KCQRL framework, these KT models do not suffer the information leakage anymore, as all models learn from the sequence of questions with our improved version.

### v2

@@ 5.2 Ablation Studies
Quality of the automated KC annotation: Here, we perform comprehensive ablations for module 1 of our framework (in Sec. 3.1) where we assess the quality of the KC annotations. (1) We ran automated evaluation of (a) the original KC annotations from the dataset and (b) the annotations from our KCQRL but without leveraging the solution steps. The evaluation is done via a different LLM, i.e., Llama-3.1-405B (Dubey et al., 2024), to avoid the potential bias from using the same model as in our KC annotation. Table 3 shows the pairwise comparison of three KC annotations (i. e., ours and two ablations) based on 5 criteria, where the LLM model is prompted to choose the best of the given two annotations for each criterion. Overall, KC annotations from our full framework are preferred over the annotations without solution steps, confirming our motivation of our KC annotation [^5] Further details and example KC annotations are in Appendix H.. (2) We ran another automated evaluation on the impact of LLM’s capacity on the quality of automated KC annotations. For this, we pick the Llama models (Dubey et al., 2024) of different sizes and used the same 5 criteria as in the earlier part. Overall, the largest LLMs consistently achieve better performance, whose details are in Appendix K. (3) After confirming the effectiveness of our KC annotation framework via automated evaluations, we finally confirm our findings via human evaluations. We we randomly selected 100 questions from the XES3G5M dataset and asked seven domain expert evaluators to compare KC annotations from three sources: (i) the original dataset, (ii) Llama-3.1-70B, and iii GPT-4o. Only 5 % of the time the evaluators preferred the original KCs over the LLM-generated KC annotations. This percentage is even lower than our automated evaluation of KC annotations in our ablation studies. The details of evaluation procedure is given in Appendix L. Takeaway: A large performance gain is due to module 1 where we ground the automated KC annotations in chain-of-thought reasoning.
Table 3: Ablation study on automated KC annotations. Preferences are shown across different criteria and datasets.
@@
Fig. 5 plots the embeddings of questions [^6] Due to space limitations, Fig. 5 shows only the embedding of (a) question content from the five ablations without any CL training. Appendix G shows the detailed visualization of all embeddings.. For all five baselines, the representations of questions from the same KC are scattered very broadly. In comparison, our KCQRL is effective in bundling the questions from the same KC in close proximity as desired. Takeaway: This highlights the importance of our CL training in Sec. 3.2.
Fig. 6 shows the performance of KT models with these five baseline embeddings in comparison to the default implementation and our complete framework. In most cases, baseline embeddings perform better than the default implementation, which again highlights the importance of semantic representations (compared to randomly initialized embeddings). Further, our KCQRL is consistently better than the baseline embeddings, which highlights the benefit of domain-specific representation learning. Takeaway: A large performance gain is from module 2, which allows our framework to capture the semantics of both questions and KCs.
Figure 5: Visualization of question embeddings. For better intuition, we chose the same question from Fig. 2, and, for each of its KCs, we color the question representations sharing the same KC. Evidently, our CL loss is highly effective.
@@
Ablations for CL training: To demonstrate the effectiveness of our CL loss, we implement the following ablations: (f) KCQRL w/o false negative elimination, whose CL loss is calculated without the blue indicator function in Eq. 3.2 and Eq. 3.2, and (g) KCQRL w/o sol. steps, whose CL loss is calculated based only on $\mathcal{L}_{\textrm{question}_{i}}$ by ignoring $\mathcal{L}_{\textrm{step}_{i}}$.
@@ 6 Related Work
Knowledge tracing (KT) models the temporal dynamics of students’ interactions with learning content to predict their knowledge over time (Corbett & Anderson, 1994). A large number of machine learning models, including deep neural networks, have been proposed for this purpose (Piech et al., 2015; Yeung & Yeung, 2018; Lee & Yeung, 2019; Sonkar et al., 2020; Long et al., 2021; Liu et al., 2023a; Chen et al., 2023; Zhang et al., 2017; Yeung, 2019; Guo et al., 2021; Pandey & Karypis, 2019; Choi et al., 2020a; Ghosh et al., 2020; Liu et al., 2023b; Huang et al., 2023).
@@
KC annotation aims to infer the KCs of a given exercise in an automated manner. (i) One research line infers KCs from students’ learning histories as latent states (e.g., Barnes, 2005; Cen et al., 2006; Liu et al., 2019; Shi et al., 2023). However, these latent states lack interpretability, limiting their use in KT. (ii) Another approach reformulates KC annotation as a classification task (e.g., Patikorn et al., 2019; Tian et al., 2022; Li et al., 2024a, b), but this requires large annotated datasets, making it often infeasible. (iii) Recent works have also explored KC annotation in free-text form (Moore et al., 2024; Didolkar et al., 2024), but their annotations are not informed by the question’s solution steps, which we show is suboptimal in our ablations. In our work, we leverage LLMs’ reasoning abilities to generate step-by-step solutions and produce grounded KCs for each question.
@@ 7 Discussion
Broader Impact
Implications: Our KCQRL allows even the simplest KT models to reach near state-of-the-art (SoTA) performance, suggesting that much of the existing literature may have “overfit” to the paradigm of modeling sequences based only on IDs. Future KT research can advance by focusing on sequential models “designed to” inherently understand semantics. As a first step, we release an English version of the XES3G5M dataset with full annotations, including question content, solutions, and KCs.
Real-world deployment: We further elaborate on how our complete KCQRL framework can be deployed and maintained in real-world scenarios. In Appendix N, we provide a comprehensive discussion on various situations that may arise when KCQRL is used for knowledge tracing in online learning platforms.
@@ Appendix A Related Work (Extended)
Knowledge tracing (KT) aims to model the temporal dynamics of students’ interactions with learning content to predict their knowledge over time (Corbett & Anderson, 1994). Early works in KT primarily leveraged logistic and probabilistic models (Corbett & Anderson, 1994; Cen et al., 2006; Pavlik et al., 2009; Käser et al., 2017; Vie & Kashima, 2019). Later, many KT works focused on deep learning-based approaches, which can loosely be grouped into different categories: 1) Deep sequential models that leverage auto-regressive architectures such as RNN or LSTM (Piech et al., 2015; Yeung & Yeung, 2018; Lee & Yeung, 2019; Liu et al., 2019; Nagatani et al., 2019; Sonkar et al., 2020; Guo et al., 2021; Long et al., 2021; Shen et al., 2022; Chen et al., 2023; Liu et al., 2023a), 2) Memory augmented models to externally capture the latent states of the students during the learning process (Zhang et al., 2017; Abdelrahman & Wang, 2019; Yeung, 2019), 3) Graph-based models that use graph neural networks to capture the interactions between knowledge concepts and questions (Nakagawa et al., 2019; Song et al., 2022; Cui et al., 2024), 4) Attention-based models that either employ a simple attention mechanism or transformer-based encoder-decoder architecture for the modeling of student interactions (Pandey & Karypis, 2019; Pandey & Srivastava, 2020; Choi et al., 2020a; Ghosh et al., 2020; Huang et al., 2023; Im et al., 2023; Liu et al., 2023b; Yin et al., 2023; Ke et al., 2024; Li et al., 2024c). Outside these categories, Lee et al. (2024) fits entire history of students exercises into the context window of a language model to make the prediction, Guo et al. (2021) designs an adversarial training to have robust representations of latent states, and Zhou et al. (2024) develops a generative model to track the knowledge states of the students. However, the existing works (1) require a comprehensive mapping between knowledge concepts and questions and (2) overlook the semantics of questions’ content and KCs. Our KCQRL framework effectively addresses both limitations.
@@
KC annotation aims at inferring the KCs of a given exercise in an automated manner. (i) One line of research infers KCs of questions by learning patterns from students’ learning histories in the form of latent states (e.g., Barnes, 2005; Cen et al., 2006; Liu et al., 2019; Shi et al., 2023). Yet, such latent states lack interpretability, because of which their use in KT is typically prohibited. (ii) Another line of research reformulates KCs annotation as a classification task, typically via model training (Patikorn et al., 2019; Tian et al., 2022) or LLM prompting (Li et al., 2024a, b). Although these works do not require human experts at inference, they still require them to curate a large annotated dataset for models to learn the task. (iii) Recent works also explored the KC annotation in free-text form (Moore et al., 2024; Didolkar et al., 2024). However, their KC annotations are not informed by the solution steps of the question. We show that this is suboptimal in our ablations.
@@
In our work, we leverage the reasoning abilities of LLMs (Wei et al., 2022; Fu et al., 2023; Shridhar et al., 2023; Wang et al., 2023; Yu et al., 2024) to generate the step-by-step solution steps of a given question and generate the associated KCs in a grounded manner.
@@
Many works designed tailored CL losses (Oord et al., 2018; Chen et al., 2020; He et al., 2020; Ozyurt et al., 2023b) to improve the representations of textual data in various domains. Examples are information retrieval (Karpukhin et al., 2020; Khattab & Zaharia, 2020; Lee et al., 2023; Sun et al., 2023a; Zhu et al., 2023; Zhao et al., 2024), textual entailment (Gao et al., 2021; Li & Li, 2024; Ou & Xu, 2024), code representation learning (Zhang et al., 2024), and question answering (Yue et al., 2021, 2022). To our knowledge, we are the first to leverage a CL loss for the representation learning of questions specific to KT tasks.
@@ Appendix B Dataset Details
XES3G5M (Liu et al., 2023c): XES3G5M contains the history student exercises for 7,652 unique questions. These questions are mapped to 865 unique KCs in total. It has a total of 18,066 students’ learning histories and a total of 5,549,635 interactions for all students. Fig. 8(a) demonstrates the distribution of interaction numbers across students. Of note, XES3G5M dataset is original provided in Chinese, which we translated to English to make it compatible to our framework. Details are given below.
@@
Translation: The original XES3G5M dataset is provided in Chinese. To make it compatible with our KCQRL framework, we first translated the question contents to English. We did the translation by Google Translate via deep-translate Python library.
@@
To make the fill-in-the-blank questions consistent with the others (and also consistent with our other dataset Eedi), We prompt GPT-4o to convert these questions to proper question phrases. From the same example, “…There are ___ different ways to wear it.” is converted to “…How many ways are there to wear it?”. We provide our prompt template in Appendix B.1.
Eedi (Eedi, 2024): Eedi contains the history student exercises for 4,019 unique questions. These questions are mapped to 1215 unique KCs in total. It has a total of 47,560 students’ learning histories and a total of 2,324,162 interactions for all students. Fig. 8(b) demonstrates the distribution of interaction numbers across students.
@@ Appendix C KT Models Details
In our KCQRL framework, we enhanced the performance of 15 KT algorithms on two real-world large online math learning datasets. The summary of these KT algorithms are given below:
  • DKT (Piech et al., 2015): It is the first deep learning based KT algorithm. It uses the LSTM to model the temporality in students’ learning histories. The original DKT turns each KC into a one-hot or a random vector. The recent implementation from pykt (Liu et al., 2022b) learns the embeddings for each KC, which are then processed by an LSTM layer.
  • DKT+ (Yeung & Yeung, 2018): It adds regularization to the existing DKT model. Specifically, it adds a reconstruction loss to the last exercise’s prediction. Further, it adds a regularization term to make the predictions of the same KC consistent across the time dimension.
@@
  • IEKT (Long et al., 2021): It models the students’ learning histories via RNN. In addition, it has two additional neural network modules, namely student cognition and knowledge acquisition estimation.
  • AT-DKT (Liu et al., 2023a): It has the same model backbone as in DKT. On top of it, AT-DKT has two auxiliary tasks, question tagging prediction and student’s individual prior knowledge prediction.
  • QIKT (Chen et al., 2023): It combines three neural network modules to model the learning processes, question-centric knowledge acquisition module via LSTM, question-agnostic knowledge state module via another LSTM and question centric problem solving module via MLP.
@@
  • simpleKT (Liu et al., 2023b): As a simple but tought to beat baseline, this model employs and ordinary dot product as an attention mechanism to the embeddings of questions.
  • sparseKT (Huang et al., 2023): On top of simpleKT, this model further introduces a sparse attention mechanism such that for the next exercise’s prediction, the model attends to at most $K$ exercises in the past.
@@ Appendix D Implementation Details
Here we explain the implementation details of each part of our KCQRL framework.
@@
KC Annotation (Sec. 3.1): We leverage the reasoning abilities of OpenAI’s GPT-4o [^7] https://platform.openai.com/docs/models/gpt-4o model as our LLM $P_{\phi}(\cdot)$ at each step. We set its temperature parameter to 0 to get the deterministic answers from the model. For each question in the dataset, GPT-4o model is prompted three times in total: one for solution steps generation, one for KC annotation and one for solution step-KC mapping. Overall, the cost is around 80 USD for XES3G5M dataset with 7,652 questions and 50 USD for Eedi dataset with 4,019 questions.
@@
Representation learning of questions Sec. 3.2: We train BERT (Devlin, 2019) as our LLM encoder $E_{\psi}(\cdot)$. For the elimination of false negative pairs, we use HDBSCAN (Campello et al., 2013) as the clustering algorithm $\mathcal{A}(\cdot)$. The clustering is done over the Sentence-BERT (Reimers & Gurevych, 2019) embeddings of KCs, which is a common practice in identifying relevant textual documents (Liu et al., 2022a; Ozyurt et al., 2023a). For HDBSCAN clustering, we set minimum cluster size to 2, minimum samples to 2, metric to cosine similarity between the embeddings. For contrastive learning training, we train BERT (Devlin, 2019) as our LLM encoder. To better distinguish three types of inputs, i. e. question content, solution step, and KC, we introduce three new tokens [Q], [S], [KC] to be learned during the training. These new tokens are added to the beginning of question content, solution step, and KC, respectively, for both training and inference. For all input types, we use [CLS] token’s embeddings as the embeddings of the entire input text. For the training objective, we set $\alpha=1$. Since both losses $\mathcal{L}_{\textrm{question}_{i}}$ and $\mathcal{L}_{\textrm{step}_{i}}$ are already normalized by the number of KCs and steps, our initial exploration ($\alpha$ varying from 0.2 to 5.0) indicated that setting $\alpha=1$ yields the best performance. We train the encoder for 50 epochs with the following hyperparameters: batch size = 32, learning rate = 5e-5, dropout = 0.1, and temperature (of similarity function) = 0.1. For training of $E_{\psi}(\cdot)$, we use Nvidia Tesla A100 with 40GB GPU memory. The entire training is completed under 4 hours.
After the training, the question embeddings are acquired by running the inference on question text and solution step texts for each question in the dataset. Specifically, we take the embeddings of "[Q] <question content>" for the question content and "[S] <solution step>" for each solution step. Then, we aggregate these embeddings as explained in Sec. 3.3. As a result, the inference does not require the KC annotations. This has the following advantage: When a new question is added to the dataset, the earlier KC annotation module can be skipped completely and one can directly get the embeddings in this module and start using them for the downstream KT model.
@@
Improving KT via learned question embeddings: We adopt the following strategy for the fair evaluation of KT algorithms and their improvements via our novel KCQRL framework. We first did a grid search over the hyperparameters of each KT model (see Table D for parameters) to find the best configuration for each model. Then, we replaced their embeddings with our learned question embeddings and trained/tested the KT models with the same configurations found earlier (i. e., no grid search is applied). To ensure the fair evaluation between the KT algorithms, and between their default versions and improved versions (via our framework), we fixed their embedding dimensions to 300. As BERT embeddings’ default dimensionality is larger (i. e., 768), we just added a linear layer (no non-linear activation function is added) on top of the replaced embeddings to reduce its dimensionality to 300 to ensure that model capacities are subjected to a fair comparison. As KT models are much smaller than the LLM encoders, this time we used NVIDIA GeForce RTX 3090 with 24GB GPU memory for even larger batch sizes.
@@
Evaluation of KC-centric KT models: Some KT models (such as DKT (Piech et al., 2015), simpleKT (Liu et al., 2023b) etc.) expand the sequence of questions into the sequence of KCs for both training and inference. For instance, if there are 3 questions with 2 KCs each, they transform the sequence $\{q_{1},\,q_{2},\,q_{3}\}$ into $\{c_{11},\,c_{12},\,c_{21},\,c_{22},\,c_{31},\,c_{32}\}$ and assign the labels of original questions to each of their corresponding KCs for training. This paradigm causes information leakage in the evaluation as highlighted by Liu et al. (2022b). The reason is, during the prediction of $c_{32}$, the label of $c_{31}$ is already given as the history, which is the same label (to be predicted) for $c_{32}$. To eliminate the information leakage, we followed the literature (Liu et al., 2022b, 2023c) and applied the following procedure. Again from the same example, to predict the label of $q_{3}$, 1) we provided the expanded KC sequence of earlier questions as before, i. e., $\{c_{11},\,c_{12},\,c_{31},\,c_{32}\}$. 2) Then we appended $c_{31}$ and $c_{32}$ separately to the given sequence and run the predictions independently. 3) Finally we aggregated these predictions by taking their mean, and used it as the model’s final prediction. Important note: With our KCQRL framework, these KT models do not suffer the information leakage anymore, as all models learn from the sequence of questions with our improved version.